OpenStudy (anonymous):

Graph the following inequality. 4x-y > 8 -

OpenStudy (anonymous):

Easiest way to graph inequalities is by getting the x and y intercept. So if y = 0 then 4x-0 > 8 4x>8 x>2 (2,0) If x = 0 then -y > 8 y < -8 (0,-8) Now, Graph both points and draw the line. Since it is only less than and not equal to, it is a dotted line and you shade downwards since when the y is positive, the sign switches.

OpenStudy (anonymous):

I get the ordered pairs easy enough but the shading part I am lost as how to graph and shade correctly, could you help me see h ow it is done?

OpenStudy (anonymous):



OpenStudy (anonymous):

Sort of like this except the line is dotted. if the sign is \[\le\] or \[\ge\], then it is a solid line. if the sign is < or > then it is a dotted line. If y is less than or less than or equal to, you shade downwards. if y is greater than or greater than or equal to, you shade upwards.

OpenStudy (anonymous):

ah, ok. Thank you for the demonstration. I understand better now. :)

OpenStudy (anonymous):

No problem ;)

Directrix (directrix):

@teenandtoddlermom In this problem, the graph should show the line to be dashed because the inequality 4x-y > 8 does not include the points on the line. As stated by @jiji501 --> if the sign is < or > then it is a dotted line.

OpenStudy (anonymous):

Yeah dotted except I was too lazy ;)